OpenStudy (anonymous):

Help PLEASE ASAP divide and simplify 3x+12/4x-18 divided by x+4/x-3 thank you!

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{3x + 12}{4x - 18} \div \dfrac{x+ 4}{x - 3} \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

\[(a/b) \div (c/d)\] can be rewritten as (a/b)*(d/c) Try this, it will help to simplify :)

OpenStudy (mathstudent55):

To divide a fraction, multiply the first fraction by the reciprocal of the second fraction. The reciprocal is simply flippipg the fraction.

OpenStudy (mathstudent55):

\(\dfrac{3x + 12}{4x - 18} \div \dfrac{x+ 4}{x - 3} \) \( =\dfrac{3x + 12}{4x - 18} \times \dfrac{x - 3}{x+ 4} \) Ok?

OpenStudy (mathstudent55):

My suggestion now is to look closely at the numerator and denominator of the left fraction. Can you factor a common factor out of the numerator? Can you factor a common factor out of the denominator?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Do so :-)